# A stochastic approach in physics exercises of mathematics education

### v1

@@ Abstract
We present a method for incorporating a stochastic point of view into physics exercises of mathematics education. The core of our method is the randomization of some inputs, the system model used does not differ from what we would use in the deterministic approach. We consider exercises from the theory of projectile motion and statics. The outputs of stochastic models are random variables, and we usually determine their probability distributions, expected values, variances, and relative standard deviations, and the probabilities of some events related to them are also calculated. We primarily recommend our exercises to motivated high school students but first-year university students can also benefit from them.
@@ 1 Introduction
In this article, we present a method for incorporating a stochastic point of view into physics exercises of mathematics education. The core of the method is the randomization of some inputs, the system model used does not differ from what we would use in the deterministic approach. In our exercises, the distribution of random inputs, in most of the cases, is a uniform distribution. This does not weaken the pedagogical goal, since using inputs with more complicated distributions would not make a difference in principle, it would only make the calculations more difficult. Of course, in real life problems, more general distributions such as exponential, normal or Pareto distributions could come into play. The outputs of stochastic models are random variables, and even if the inputs are uniformly distributed, the outputs are usually not uniformly distributed. We note that in Exercise 2.4 an absolutely continuous random variable with a given density function also appears for the ratio of the rebound and landing speeds when a tennis ball is dropped on concrete from a given distance. In the exercises, we usually determine the probability distributions, expected values, variances, and relative standard deviations of the outputs, and the probabilities of some events related to them are also calculated using the basic tools of probability theory.
@@
Our reason for choosing the topics of projectile motion and statics for demonstrating our method is the fact that the physics behind these two topics is relatively easy to understand, and we would like to recommend our exercises to motivated high school students. Students learn the physics of projectile motion as early as the first years of high school. However, first-year university students can also benefit from getting acquainted with these exercises. Students may be asked to create similar exercises, for other fields of physics as well.
In Appendices A and B, we collected some of the basics of probability theory and mechanics that are needed to solve the problems, respectively. Most of the content of Appendix A on the required knowledge of probability theory is covered by the syllabus of high school students specialized in mathematics, except maybe the law of total expectation. Nonetheless, we remark that the law of total expectation is used only in the solution of part (ii) of Exercise 2.1 where an alternative solution using the method of geometric probability is also presented. The content of Appendix B on some knowledge of mechanics is essentially contained in the syllabus of high school students, and hence this appendix is not so detailed.
@@
We take great care to demonstrate the use of the tools from probability theory, and hence we provide detailed solutions to the exercises. Our stochastic perspective can help the students recognize the fact that many phenomena in our daily lives have a random nature, and therefore it can be useful to get in touch with stochastic thinking and tools in order to be able to understand these phenomena much better.
@@
The paper is structured as follows. Sections 2 and 3 are devoted to exercises for projectile motion and statics, respectively. In Sections 2 and 3, we present five and three exercises, respectively. We close the paper with Appendices A and B, where, without being exhaustive, we recall some of the notations, conventions, concepts and results of probability theory and mechanics that are used in solving the exercises.
@@
Next, we briefly summarize what kind of tools in probability theory we use in the exercises of Sections 2 and 3. In Exercise 2.1, among others, we use the notions of expectation, variance and relative standard deviation of an absolutely continuous random variable, and we calculate the probability of an event by two methods, namely, by the method of geometric probability and by the law of total expectation, respectively. In part (iii) of Exercise 2.1, we calculate a conditional probability as well. In Exercise 2.2, we use the notions of expectation, variance and relative standard deviation of a discrete random variable, and the additivity property of probability. In Exercise 2.3, among others, the notion of the distribution of discrete and absolute continuous random variables and the notions of expectations, variances and relative standard deviations of such random variables come into play. In Exercise 2.4, we use the fact that a probability that an absolutely continuous random variable falls in a given interval can be calculated using the density function of the random variable in question, and we provide an application of Bayes’ theorem as well. In Exercise 2.5, we calculate the expected value of the relative error in some problems for free fall (i.e., for a motion of a body, where the gravity of Earth is the only force acting upon it) if instead of the precise value of the gravity of Earth we use an approximating random value for it. In this exercise, the notion of the expected value of an absolutely continuous random variable is used. In Exercises 3.2 and 3.3, 2-dimensional random variables and their expectation vectors play important roles.
The set of positive integers and real numbers is denoted by $\mathbb{N}$ and $\mathbb{R}$, respectively. For a $2$-dimensional vector $(x,y)\in\mathbb{R}^{2}$, its Euclidean length is defined by $\|(x,y)\|:=\sqrt{x^{2}+y^{2}}$. The coordinates of a vector ${\boldsymbol{v}}$ in the $(x,y)$-plane are denoted by $v_{x}$ and $v_{y}$, respectively, i.e., ${\boldsymbol{v}}=(v_{x},v_{y})$. In the solutions, the number of significant digits is chosen to be 4 (for example, instead of $21.285$, we write $21.29$, or instead of $0.023456$, we write $0.02346$). All of the physical quantities are expressed in SI units, so we do not display the units in calculations. In the calculations, the magnitude of the gravitational acceleration near the Earth’s surface is assumed to be $g=9.81\,$$\mathrm{\frac{m}{s^{2}}}$.
@@ 2 A stochastic approach in projectile motion
In this section, we demonstrate our method for incorporating a stochastic point of view into physics exercises of mathematics education by presenting exercises for projectile motion. In these ones, some of the initial values, namely, the initial height and/or velocity of a launched ball, are chosen randomly to make the problems more realistic. We usually choose a uniform distribution, but in Exercise 2.4 an absolutely continuous random variable with a given density function also comes into play for the ratio of the rebound and landing speeds when a tennis ball is dropped on concrete from a given distance. In all of the exercises, we suppose that the air resistance is negligible. Altogether, we present five problems in this section.
@@
Let us consider a cylindrical-shape forest lookout with a radius of 10 meters and a height of 50 meters, which stands on horizontal ground. From the bottom of the forest lookout, Thomas goes up the spiral staircase. Thomas stops at a height $h$ chosen according to a uniform distribution on the interval $[0\,\mathrm{m},50\,\mathrm{m}]$, and from the edge of the forest lookout he launches a ball with velocity $v$ of which the magnitude is uniformly distributed on the interval $[5\,\mathrm{\frac{m}{s}},15\,\mathrm{\frac{m}{s}}]$, and its direction is perpendicular to the wall of the forest lookout. We suppose that $h$ and $v$ are independent.
@@
    Suppose that the forest lookout is surrounded by an annulus-shape fence to which the top is horizontal. The two walls of the fence (inner and outer circles) are $10$ meters and $10.5$ meters from the wall of the forest lookout, respectively. The height of the fence is $2$ meters. On Figure 1, the plane section determined by the direction of the shot is shown, and only the part of the fence in the direction of the shot is plotted. What is the probability that the ball first hits one of the line segments $BC$ and $CD$ (see Figure 1), that is, the ball first hits the fence?
    Figure 1:
@@
Figure 2:
@@
$R:=\Big{\{}\text{the ball's (first) point of impact is in the line segment $AB$}\Big{\}}=\Big{\{}d<10\Big{\}},$
$S:=\Big{\{}\text{the ball's (first) point of impact is to the right of the point $E$}\Big{\}}$
=$\phantom{:=\,}=\Big{\{}\frac{g}{2v^{2}}\cdot 10.5^{2}<h-2\Big{\}}.$
@@
(2.1) ({the ball hits the fence before it hits the $\operatorname{\mathbb{P}}\big{(}\text{\{the ball hits the fence before it hits the ground\}}\big{)}=1-\operatorname{\mathbb{P}}(R)-\operatorname{\mathbb{P}}(S).$
In what follows, we determine the probabilities $\operatorname{\mathbb{P}}(R)$ and $\operatorname{\mathbb{P}}(S)$ using two different methods: the method of geometric probabilities (Method 1) and the law of total expectation (Method 2).
Method 1 (geometric probabilities): By Figure 3
Figure 3:
@@
where $500$ is the area of the rectangle $[0,50]\times[5,15]$ (”total” area), and the ”probable” area $T_{R}$ can be calculated as follows. First, we determine the points at which the function $(0,\infty)\ni h\mapsto\frac{10\sqrt{g}}{\sqrt{2}}\cdot\frac{1}{\sqrt{h}}$ takes the values $15$ and $5$, respectively (see Figure 3). By algebraic calculations, we can get that
@@
Figure 4:
@@
Method 2 (law of total expectation): Since $v$ is uniformly distributed on the interval $[5,15]$, its density function takes the form
where
since $19.6250g/r^{2}\leqslant 50g/25=$ 19.62. Consequently,
Similarly, using the law of total expectation (see (A.4)) and the independence of $h$ and $v$, we have that
Note that we calculated the same value for $\operatorname{\mathbb{P}}(S)$ using Methods 1 and 2, as it was expected.
@@
$1-\operatorname{\mathbb{P}}(R)-\operatorname{\mathbb{P}}(S)=1-\frac{g}{75}-\frac{48}{50}+\frac{110.25g}{7500}=\frac{2}{50}+\left(\frac{110.25}{7500}-\frac{1}{75}\right)g\approx 0.05341$
@@
Figure 5:
@@
where $500$ is the area of the rectangle $[0,50]\times[5,15]$ (”total” area), and the ”probable” area $T$ can be calculated as follows. Using (2.2) and (2.3), we have that
@@
We note that students may formulate and solve a natural counterpart of Exercise 2.1 by choosing the height (where Thomas stops) according to a (discrete) uniform distribution, for example, on the finite set $\{1\,\mathrm{m},2\,\mathrm{m},\ldots,49\,\mathrm{m},50\,\mathrm{m}\}$. Of course, one could also consider other discrete or absolutely continuous distributions with range in $[0,50]$. For example, the discrete distribution on the set $\{10\,\mathrm{m},20\,\mathrm{m},30\,\mathrm{m},40\,\mathrm{m},50\,\mathrm{m}\}$ with corresponding probabilities $\frac{40}{100},\frac{30}{100},\frac{20}{100},\frac{9}{100},\frac{1}{100}$ would may suggest that Thomas has a fear of heights.
@@
Part (i) of the next Exercise 2.3 is a randomized version of Exercise 2 for 4-th grade high school students at the Ágoston Budó Physics Competition (1980) (see [6]) in the sense that we randomly choose the height from which we drop the ball. Parts (ii), (iii) and (iv) of Exercise 2.3 contain new probabilistic-type questions to the same phenomena.
@@
From a height chosen uniformly distributed between 00 and $1$ meters, we drop a ball on the table. We assume that the ball loses $5\%$ of its velocity by each collision with the table.
@@
(iv). The range of the random variable $\xi$ is $\{0,1,2,\ldots\}$, but we will check that $\xi$ takes every positive even number with probability 00, and it will also turn out that there exists a natural number $n_{0}$ such that the probability that $\xi$ takes a value greater than $n_{0}$ is 00.
@@
$\operatorname{\mathbb{P}}(\xi=2k+1)$ =($=\operatorname{\mathbb{P}}\left(\{h_{0}>0.1\}\cap\bigcap_{j=1}^{k}\Big{\{}\frac{v_{j}^{2}}{2g}>0.1\Big{\}}\cap\Big{\{}\frac{v_{k+1}^{2}}{2g}<0.1\Big{\}}\right)$
=($=\operatorname{\mathbb{P}}\left(\{h_{0}>0.1\}\cap\bigcap_{j=1}^{k}\Big{\{}\frac{(0.95^{j}\cdot v_{0})^{2}}{2g}>0.1\Big{\}}\cap\Big{\{}\frac{(0.95^{k+1}\cdot v_{0})^{2}}{2g}<0.1\Big{\}}\right)$
=($=\operatorname{\mathbb{P}}\left(\{h_{0}>0.1\}\cap\bigcap_{j=1}^{k}\Big{\{}\frac{0.95^{2j}\cdot 2gh_{0}}{2g}>0.1\Big{\}}\cap\Big{\{}\frac{0.95^{2(k+1)}\cdot 2gh_{0}}{2g}<0.1\Big{\}}\right)$
@@
=$=\begin{cases}\frac{0.1}{0.95^{2k}}\left(\frac{1}{0.95^{2}}-1\right)&\text{if \ $\frac{0.1}{0.95^{2(k+1)}}<1$,}\\ 1-\frac{0.1}{0.95^{2k}}&\text{if \ $\frac{0.1}{0.95^{2k}}<1<\frac{0.1}{0.95^{2(k+1)}}$,}\\ 0&\text{if \ $\frac{0.1}{0.95^{2k}}>1$.}\end{cases}$
@@
Note that the function $\{0,1,\ldots,21\}\ni k\mapsto\operatorname{\mathbb{P}}(\xi=2k+1)$ is strictly increasing, however, the function $\{0,1,\ldots,22\}\ni k\mapsto\operatorname{\mathbb{P}}(\xi=2k+1)$ is not strictly increasing, since $\operatorname{\mathbb{P}}(\xi=43)\approx 0.09315$, which is greater than $\operatorname{\mathbb{P}}(\xi=45)\approx 0.04465$. Remark also that the (unique) maximum of the probabilities $\operatorname{\mathbb{P}}(\xi=0)$ and $\operatorname{\mathbb{P}}(\xi=2k+1)$, $k\in\{0,1,\ldots,22\}$ is equal to $\operatorname{\mathbb{P}}(\xi=0)=0.1$, yielding that the mode of the distribution of $\xi$ is 00. $\Box$
@@
i.e., a tennis ball produced by Company B rebounds to the expected height of approximately 1.494 cm dropped from a height of 254 cm.
@@
{a tennis ball is $\Big{\{}\text{a tennis ball is standard}\Big{\}}=\{h_{f}\in(1.35,1.47)\},$
@@
Part (i) of the next Exercise 2.5 is a randomized version of Exercise 1, Round II, Category III at the Sándor Mikola Physics Competition (2023) (see [4]) in the sense that the precise theoretical value of the gravity of Earth is considered to be a random value. Note that this kind of normalization is different from the ones in Exercises 2.1-2.3 in the sense that the uncertainty in a system’s parameter is studied in Exercise 2.5 instead of the uncertainty in some of the inputs of a dynamical system.
@@
Let $g$ denote the (theoretical) value of the gravity of Earth and let $\xi$ be a uniformly distributed random variable on the interval $[-1,1]$. What is the expectation of the relative error that we make if the gravity of Earth is taken to be $(g+\xi)$ instead of the theoretical value $g$ and
@@
that is, the expected value of the relative error in question is 00.
@@
This expected value is positive, since the function $[0,\infty)\ni x\mapsto\sqrt{x}$ is strictly concave, and hence

### v2

@@ Abstract
We present a method for incorporating a stochastic point of view into physics exercises of mathematics education. The core of our method is the randomization of some inputs, the system model used does not differ from what we would use in the deterministic approach. We consider exercises from the theory of projectile motion and statics. The outputs of stochastic models are random variables, and we usually determine their probability distributions, expected values, variances, and relative standard deviations, and the probabilities of some events related to them are also calculated. Students and teachers familiar with elementary probability theory and mechanics may find these exercises useful for understanding some basic concepts of stochastic mechanics.
@@ 1 Introduction
In this article, we present a method for incorporating a stochastic point of view into physics exercises of mathematics education. The core of the method is the randomization of some inputs, the system model used does not differ from what we would use in the deterministic approach. We take as a basis the mathematical models of some real physical phenomena and we incorporate randomness into this mathematical model by randomizing some of the inputs. In our exercises, the distribution of random inputs, in most of the cases, is a uniform distribution. This does not weaken the pedagogical goal, since using inputs with more complicated distributions would not make a difference in principle, it would only make the calculations more difficult. Of course, in real life problems, where the distribution of the random inputs can be more general than uniform distributions, as a first step, one should approximate the distribution functions of the inputs, for example, by the help of empirical distribution functions. The outputs of stochastic models are random variables, and even if the inputs are uniformly distributed, the outputs are usually not uniformly distributed. We note that in Exercise 2.5 an absolutely continuous random variable with a given density function also appears for the ratio of the rebound and landing speeds when a tennis ball is dropped on concrete from a given distance. In the exercises, we usually determine the probability distributions, expected values, variances, and relative standard deviations of the outputs, and the probabilities of some events related to them are also calculated using the basic tools of probability theory. In real life problems, one can usually give only approximations of the quantities in question (possibly using some software), since, as we mentioned above, in practice the distributions of the random inputs are approximated as well. We emphasize that we do not develop any new mathematical model for physical phenomena, we just incorporate randomness into the existing models by considering random inputs having uniform distributions in most of the cases (for simplicity). Having stochastic models at hand, in our exercises, we present several methods for solving probabilistic problems related to the random outputs of the models in question.
@@
Our reason for choosing the topics of projectile motion and statics for demonstrating our method is the fact that the physics behind these two topics is relatively easy to understand. In Appendices A and B, we collected some of the basics of probability theory and mechanics that are needed to solve the problems, respectively. We mention that, in Appendix A and in the solutions of some of the exercises, absolutely continuous random variables and their expected value and variance come into play. In high schools it may be not part of the standard curriculum, but in such a situation interested teachers can include it for the sake of interested students. At universities, every course on basic probability theory contains the topic of absolutely continuous random variables as well. The content of Appendix B on some knowledge of mechanics is essentially contained in the syllabus of high school students, and hence this appendix is not so detailed.
The solutions of our exercises can be successfully understood by anyone who is familiar with the content of Appendices A and B. In particular, by high school students who undertake extension work in the mathematical sciences. In some classes in specialized mathematics, high school students can have 7 or 8 hours per week in mathematics, and the teachers may have the possibility to present material from probability theory that are even not included in the ”standard” syllabus (such as absolutely continuous random variables). In this way, students can also practice the integration of some elementary functions that is part of the ”standard” syllabus in specializes math classes. Further, there are so-called non-compulsory workshops, study circles for interested math students, who may prepare for some competitions, and at these occasions the teachers may also consider our randomized exercises. Concerning university students, we think that our randomized exercises are definitely understandable for those university students who took a course on basic probability theory.
@@
We take great care to demonstrate the use of the tools from probability theory, and hence we provide detailed solutions to the exercises. Our stochastic perspective can help the students recognize the fact that many phenomena in our daily lives have a random nature, and therefore it can be useful to get in touch with stochastic thinking and tools in order to be able to understand these phenomena much better. Our exercises are related to projectile motion and statics, and we mention that STEM students may be asked to create similar exercises, for other fields of physics as well.
@@
The paper is structured as follows. Sections 2 and 3 are devoted to exercises for projectile motion and statics, respectively. In Sections 2 and 3, we present five and three exercises, respectively. We close the paper with three appendixes. In Appendices A and B, without being exhaustive, we recall some of the notations, conventions, concepts and results of probability theory and mechanics that are used in solving the exercises. Appendix C is devoted to an alternative solution for calculating some probabilities for part (ii) of Exercise 2.1 using the law of total expectation instead of the method of geometric probabilities.
@@
Next, we briefly summarize what kind of tools in probability theory we use in the exercises of Sections 2 and 3. In Exercise 2.1, among others, we use the notions of expectation, variance and relative standard deviation of an absolutely continuous random variable, and we calculate the probability of an event by two methods, namely, by the method of geometric probability and by the law of total expectation, respectively. In part (iii) of Exercise 2.1, we calculate a conditional probability as well. In Exercise 2.3, we use the notions of expectation, variance and relative standard deviation of a discrete random variable, and the additivity property of probability. In Exercise 2.4, among others, the notion of the distribution of discrete and absolute continuous random variables and the notions of expectations, variances and relative standard deviations of such random variables come into play. In Exercise 2.5, we use the fact that a probability that an absolutely continuous random variable falls in a given interval can be calculated using the density function of the random variable in question, and we provide an application of Bayes’ theorem as well. In Exercise 2.6, we calculate the expected value of the relative error in some problems for free fall (i.e., for a motion of a body, where the gravity of Earth is the only force acting upon it) if instead of the precise value of the gravity of Earth we use an approximating random value for it. In this exercise, the notion of the expected value of an absolutely continuous random variable is used. In Exercises 3.2 and 3.3, 2-dimensional random variables and their expectation vectors play important roles.
[…]
@@ 2 A stochastic approach in projectile motion
In this section, we demonstrate our method for incorporating a stochastic point of view into physics exercises of mathematics education by presenting exercises for projectile motion. In these ones, some of the initial values, namely, the initial height and/or velocity of a launched ball, are chosen randomly to make the problems more realistic. We usually choose a uniform distribution, but in Exercise 2.5 an absolutely continuous random variable with a given density function also comes into play for the ratio of the rebound and landing speeds when a tennis ball is dropped on concrete from a given distance. In all of the exercises, we suppose that the air resistance is negligible. Altogether, we present five problems in this section. These exercises introduce a simple application of probability theory to mechanical problems, they present composite systems influenced by random parameters. Students can see how variability in measurements can be modeled on a probabilistic basis.
@@
Let us consider a cylindrical-shape forest lookout with a radius of 10 meters and a height of 50 meters, which stands on horizontal ground. From the bottom of the forest lookout, Thomas goes up the spiral staircase. Thomas stops at a height $h$ chosen according to a uniform distribution on the interval $0\,\mathrm{m},50\,\mathrm{m}]$, and from the edge of the forest lookout he launches a ball with velocity $v$ of which the magnitude is uniformly distributed on the interval $5\,\mathrm{\frac{m}{s}},15\,\mathrm{\frac{m}{s}}]$, and its direction is perpendicular to the wall of the forest lookout. We suppose that $h$ and $v$ are independent.
@@
    Suppose that the forest lookout is surrounded by an annulus-shape fence to which the top is horizontal. The two walls of the fence (inner and outer circles) are $10$ meters and $10.5$ meters from the wall of the forest lookout, respectively. The height of the fence is $2$ meters. On Figure 1, the planar section determined by the direction of the shot is shown, and only the part of the fence in the direction of the shot is plotted. What is the probability that the ball first hits one of the line segments $BC$ and $CD$ (see Figure 1), that is, the ball first hits the fence?
    Figure 1: a planar section of the forest lookout surrounded by an annulus-shape fence for parts (ii) and (iii) in Exercise 2.1.
@@
Figure 2: the trajectory of the motion of a ball launched horizontally at height $h$ with velocity $v$ for the solution of Exercise 2.1.
@@
[…]
$\phantom{:=\,}=\Big{\{}\frac{g}{2v^{2}}\cdot 10.5^{2}<h-2\Big{\}}.$
@@
[…]
In what follows, we determine the probabilities $\operatorname{\mathbb{P}}(R)$ and $\operatorname{\mathbb{P}}(S)$ using the method of geometric probabilities. In Appendix C, as an alternative solution, we determine these two probabilities using the law of total expectation as well.
By Figure 3
Figure 3: for calculating the probability $\operatorname{\mathbb{P}}(R)$ in the solution of part (ii) in Exercise 2.1.
@@
where $500$ is the area of the rectangle $0,50]\times[5,15]$ (”total” area), and the ”probable” area $T_{R}$ can be calculated as follows. First, we determine the points at which the function $(0,\infty)\ni h\mapsto\frac{10\sqrt{g}}{\sqrt{2}}\cdot\frac{1}{\sqrt{h}}$ takes the values $15$ and $5$, respectively (see Figure 3). By algebraic calculations, we can get that
@@
Figure 4: for calculating the probability $\operatorname{\mathbb{P}}(S)$ in the solution of part (ii) in Exercise 2.1.
@@
$1-\operatorname{\mathbb{P}}(R)-\operatorname{\mathbb{P}}(S)=1-\frac{g}{75}-\frac{48}{50}+\frac{110.25g}{7500}=\frac{2}{50}+\left(\frac{110.25}{7500}-\frac{1}{75}\right)g\approx 0.05341.$
@@
Figure 5: for calculating the probability $\operatorname{\mathbb{P}}(h<10,d<10)$ in the solution of part (iii) in Exercise 2.1.
@@
where $500$ is the area of the rectangle $0,50]\times[5,15]$ (”total” area), and the ”probable” area $T$ can be calculated as follows. Using (2.2) and (2.3), we have that
@@
2.2 Remark.
(i). We compare the solution of part (i) of Exercise 2.1 with the case when the height (where Thomas stops) is chosen to be $25\,\mathrm{m}$ and the magnitude of the velocity of the ball is chosen to be $10\,\mathrm{\frac{m}{s}}$. Note that the values $25$ and $10$ are nothing else but the expected values of the corresponding uniformly distributed random variables considered in Exercise 2.1 for the height and magnitude of velocity, respectively. Then the total (horizontal) distance of the whole journey before the ball hits the ground is $\sqrt{\frac{2\cdot 25}{g}}\cdot 10\approx 22.58$. We call the attention to the fact that $\operatorname{\mathbb{E}}(d)$ calculated in the solution of part (i) in Exercise 2.1 is not equal to $\sqrt{\frac{50}{g}}\cdot 10$. The relative error of the two quantities in question is
$\frac{\operatorname{\mathbb{E}}(d)}{\sqrt{\frac{50}{g}}\cdot 10}=\frac{\frac{200}{3\sqrt{g}}}{\sqrt{\frac{50}{g}}\cdot 10}=\frac{2\sqrt{2}}{3}<1.$
(ii). We note that students may formulate and solve some natural counterparts of Exercise 2.1. Namely, one could choose the height (where Thomas stops) according to a (discrete) uniform distribution, for example, on the finite set $\{1\,\mathrm{m},2\,\mathrm{m},\ldots,49\,\mathrm{m},50\,\mathrm{m}\}$. Of course, one could also consider other discrete or absolutely continuous distributions with range in $[0,50]$. For example, the discrete distribution on the set $\{10\,\mathrm{m},20\,\mathrm{m},30\,\mathrm{m},40\,\mathrm{m},50\,\mathrm{m}\}$ with corresponding probabilities $\frac{40}{100},\frac{30}{100},\frac{20}{100},\frac{9}{100},\frac{1}{100}$ would may suggest that Thomas has a fear of heights. As for another variant of Exercise 2.1, one could consider a case when only one of the inputs is randomized, for example, one could choose the height (where Thomas stops) according to some probability distribution, but the magnitude of the velocity of the ball is taken to be a fixed (non-random) value. $\Box$
@@
Part (i) of the next Exercise 2.4 is a randomized version of Exercise 2 for 4-th grade high school students at the Ágoston Budó Physics Competition (1980) (see [6]) in the sense that we randomly choose the height from which we drop the ball. Parts (ii), (iii) and (iv) of Exercise 2.4 contain new probabilistic-type questions to the same phenomena.
@@
From a height chosen uniformly distributed between 0 and $1$ meters, we drop a ball on the table. We assume that the ball loses $5\%$ of its velocity by each collision with the table.
@@
(iv). The range of the random variable $\xi$ is $\{0,1,2,\ldots\}$, but we will check that $\xi$ takes every positive even number with probability 0, and it will also turn out that there exists a natural number $n_{0}$ such that the probability that $\xi$ takes a value greater than $n_{0}$ is 0.
@@
$\operatorname{\mathbb{P}}(\xi=2k+1)$ $=\operatorname{\mathbb{P}}\left(\{h_{0}>0.1\}\cap\bigcap_{j=1}^{k}\Big{\{}\frac{v_{j}^{2}}{2g}>0.1\Big{\}}\cap\Big{\{}\frac{v_{k+1}^{2}}{2g}<0.1\Big{\}}\right)$
$=\operatorname{\mathbb{P}}\left(\{h_{0}>0.1\}\cap\bigcap_{j=1}^{k}\Big{\{}\frac{(0.95^{j}\cdot v_{0})^{2}}{2g}>0.1\Big{\}}\cap\Big{\{}\frac{(0.95^{k+1}\cdot v_{0})^{2}}{2g}<0.1\Big{\}}\right)$
$=\operatorname{\mathbb{P}}\left(\{h_{0}>0.1\}\cap\bigcap_{j=1}^{k}\Big{\{}\frac{0.95^{2j}\cdot 2gh_{0}}{2g}>0.1\Big{\}}\cap\Big{\{}\frac{0.95^{2(k+1)}\cdot 2gh_{0}}{2g}<0.1\Big{\}}\right)$
@@
$=\begin{cases}\frac{0.1}{0.95^{2k}}\left(\frac{1}{0.95^{2}}-1\right)&\text{if \ $\frac{0.1}{0.95^{2(k+1)}}<1$,}\\ 1-\frac{0.1}{0.95^{2k}}&\text{if \ $\frac{0.1}{0.95^{2k}}<1<\frac{0.1}{0.95^{2(k+1)}}$,}\\ 0&\text{if \ $\frac{0.1}{0.95^{2k}}>1$.}\end{cases}$
@@
Note that the function $\{0,1,\ldots,21\}\ni k\mapsto\operatorname{\mathbb{P}}(\xi=2k+1)$ is strictly increasing, however, the function $\{0,1,\ldots,22\}\ni k\mapsto\operatorname{\mathbb{P}}(\xi=2k+1)$ is not strictly increasing, since $\operatorname{\mathbb{P}}(\xi=43)\approx 0.09315$, which is greater than $\operatorname{\mathbb{P}}(\xi=45)\approx 0.04465$. Remark also that the (unique) maximum of the probabilities $\operatorname{\mathbb{P}}(\xi=0)$ and $\operatorname{\mathbb{P}}(\xi=2k+1)$, $k\in\{0,1,\ldots,22\}$ is equal to $\operatorname{\mathbb{P}}(\xi=0)=0.1$, yielding that the mode of the distribution of $\xi$ is 0. $\Box$
@@
i.e., a tennis ball produced by Company B rebounds to the expected height of approximately 149.4 cm dropped from a height of 254 cm.
@@
$\Big{\{}\text{a tennis ball is standard}\Big{\}}=\{h_{f}\in(1.35,1.47)\},$
@@
Part (i) of the next Exercise 2.6 is a randomized version of Exercise 1, Round II, Category III at the Sándor Mikola Physics Competition (2023) (see [4]) in the sense that the precise theoretical value of the gravity of Earth is considered to be a random value. Note that this kind of normalization is different from the ones in Exercises 2.1-2.4 in the sense that the uncertainty in a system’s parameter is studied in Exercise 2.6 instead of the uncertainty in some of the inputs of a dynamical system.
@@
Let $g$ denote the (theoretical) value of the gravity of Earth and let $\xi$ be a uniformly distributed random variable on the interval [1,1][-1,1]. What is the expectation of the relative error that we make if the gravity of Earth is taken to be $(g+\xi)$ instead of the theoretical value $g$ and
@@
that is, the expected value of the relative error in question is 0.
@@
This expected value is positive, since the function $0,\infty)\ni x\mapsto\sqrt{x}$ is strictly concave, and hence
@@
Regarding didactic conclusions, we can say that our exercises are suitable for introducing some basic concepts of probability theory, such as expected value, variance and density function of a random variable, in a physical context. They enable comparison between stochastic and deterministic solutions (for example, in case of part (i) of Exercise 2.1, see part (i) of Remark 2.2). Exercise 2.5 encourages the exploration of different probability distributions to describe the behavior of a variable system. Moreover, students can extend the method to similar systems to develop their problem-solving skills (for example, for some natural counterparts of Exercise 2.1, see part (ii) of Remark 2.2).